OpenStudy (anonymous):

x(2x+3) distributive property

OpenStudy (anonymous):

x*2x=3x x*3=3x 3x+3x=6x

OpenStudy (anonymous):

no because the 2x becomes squared correct?

OpenStudy (anonymous):

no. because broken down it would be x*x*x. how many x's are there?

OpenStudy (anonymous):

no. because broken down it would be x*x*x. how many x's are there?

OpenStudy (anonymous):

well it is bacsically 2 multiplied by x but x squared is x *x

OpenStudy (anonymous):

x*2x + 3x = 2x^2 +3x

OpenStudy (anonymous):

uggg know i have to make it equal 150

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

sorry pudgesan, tejasree is right. by friend had me messed up!

OpenStudy (anonymous):

thats cool good try

OpenStudy (anonymous):

now you have to make that equation equal 150?

OpenStudy (anonymous):

yeah problem solving the length of a rectangle is 3 more com that twice the width the perimeter is 150 and i have to find the length

OpenStudy (anonymous):

so its.... 2x^2 + 3x = 150?

OpenStudy (anonymous):

wait thats wrong

OpenStudy (anonymous):

no its not

OpenStudy (anonymous):

because you are asking for perimeter, the one that I gave you is for area

OpenStudy (anonymous):

the width is x and the length is 2x +3, right?

OpenStudy (anonymous):

oh so how do i find the perimeter

OpenStudy (anonymous):

P = 2l + 2w 150 = 2(2x+3) + 2(x) 150 = 4x + 6 + 2x 150 = 6x +6 150-6 = 6x + 6 - 6 144 = 6x 24 = x

OpenStudy (anonymous):

your amazing thank you so much

OpenStudy (anonymous):

no its ok, thank you! Oh thanks for becoming my fan and giving me my first medal!!YAY!!!!!

OpenStudy (anonymous):

welcome thanks for saving my butt

OpenStudy (anonymous):

lol! :)